In the computer language Python, indicators must start with an upper-case letter, a lower-case

Can you explain your answer and logic step by steps. and state any laws or def. you used Thank you!
Answer #1

Let's consider 8 unfilled positions that we want to fill with characters(comprising of Uppercase letters, lowercase letters, digits and underscore) such that they form identifiers.

_ _ _ _ _ _ _ _ (8 unfilled positions)

Now, to fill the first position, we have 53 characters(26 Uppercase letters + 26 Lowercase letters + 1 underscore).

Therefore, Number of Identifiers with length 1 = 53.

Now, to fill the second position, we have 63 characters(26 Uppercase letters+26 Lowercase letters + 1 Underscore + 10 digits[0-9]).

Therefore, we have 63 characters to fill the second position.

Since an identifier with length 2 will be combination of characters at position 1 and position 2,

the number of identifiers with length 2 = 53(for first position) X 63(for second position) = 53 * 63 = 3339

Similarly, for third position, we have 63 characters.

Number of identifiers with length 3 = 53 * 63 * 63 = 53 * 632

Number of identifiers with length 4 = 53 * 63 * 63 * 63 = 53 * 633

Number of identifiers with length 5 = 53 * 63 * 63 * 63 * 63= 53 * 634

Number of identifiers with length 6 = 53 * 63 * 63 * 63 * 63 * 63= 53 * 635

Number of identifiers with length 7 = 53 * 63 * 63 * 63 * 63 * 63 * 63= 53 * 636

Number of identifiers with length 8 = 53 * 63 * 63 * 63 * 63 * 63 * 63 * 63= 53 * 637

Sum of these will give the total number of words with length <=8 but these will contain the reserved words. We have to subtract the number of reserved words to get the total number of identifiers.

Therefore, number of identifiers with length <=8 = 53 + 53*63 +  53 * 632 + 53 * 633 + 53 * 634 + 53 * 635+ 53 * 636 + 53 * 637 - 29 = 2.1213317 X 1014 . .

Hope you get it.